- A muni is offering a yield of 3.5%. Your firm has a 25% marginal tax rate. What’s the muni’s TEY?

   - A T-bill has a face value of $100,000. It has a maturity of 90 days and can be purchased at a discount rate of 1%.   The purchase price is $99,000 ($100,000 − $1,000).

What is the T-bill’s money market yield (MMY) and bond-equivalent yield (BEY)?
Answer #1

tax equivalent yield of Munis =Yield/(1-Tax Rate)= 3.5%/(1-25%) =4.67%.

MMY of T-bill =1000/(100000-1000)*(360/90)=4.04%
BEY of t-bill=1000/(100000-1000)*(360/95)=4.10%